please could you help with rearranging this equation.....
x=7+8y / 7+8y Express y in terms of x. thank you!!!!

Answer by CharlesG2(834)

please could you help with rearranging this equation.....
x=7+8y / 7+8y Express y in terms of x. thank you!!!!


if problem is x = (7 + 8y)/(7 + 8y), then:
x = (7 + 8y)/(7 + 8y)
x(7 + 8y) = 7 + 8y (multiplied both sides by 7 + 8y)
7x + 8xy = 7 + 8y
8xy - 8y = 7 - 7x (subtracted 7x from both sides, subtracted 8y from both sides)
y(8x - 8) = 7 - 7x
y = (7 - 7x)/(8x - 8)


if this was instead the problem: x = 7 + 8y/7 + 8y, then:
x = 7 + 8y/7 + 8y
-8y/7 - 8y = 7 - x (subtracted 8y/7 + 8y from both sides, subtracted x from both sides)
-8y - 56y = 49 - 7x (multiplied both sides by 7)
-64y = 49 - 7x
64y = 7x - 49 (divided both sides by -1)
y = (7x - 49)/64